OpenStudy (anonymous):

Let f(x) = 9x - 2 and g(x) = -x + 3. Find f(g(x)).

OpenStudy (anonymous):

f(g(x)) = 9*(-x+3) - 2 = -9x + 27 - 2 f(g(x)) = -9x +25